OpenStudy (ayyookyndall):

Geometry: How does the length of the hypotenuse in a right triangle relate to the lengths of the legs?

OpenStudy (johnweldon1993):

Via the pythagoran theorem \[\large a^2 + b^2 = c^2\] c is the hypotenuse of the right triangle and a and b are the legs

OpenStudy (anonymous):

Look up Pythagoras' Theorem and see if you can tell me the answer.

OpenStudy (ayyookyndall):

Can you help me with one more? @johnweldon1993

OpenStudy (johnweldon1993):

There are many people who can so if I cannot then I'm sure someone else will :)

OpenStudy (ayyookyndall):

Okay. How do you find the value of x? What is it?

OpenStudy (johnweldon1993):

We have 2 legs known...and 1 angle...if we use the law of cosines to solve for side c...and then the law of sines to solve for the missing angle \[\large c^2 = a^2 + b^2 - 2ab(cos(c))\] Doing that...we can now see that we have 3 sides...and we set up the law of sines \[\large \frac{sin(a)}{A} = \frac{sin(b)}{B} = \frac{sin(c)}{C}\] So we plug in whatever sides we know..and solve for the angle 'x' \[\large \frac{sin(38)}{\text{side c we just solved for with the law of cosines}}=\frac{sin(x)}{21}\] and solve for 'x'

OpenStudy (zarkon):

don't work so hard...this is an isosceles triangle

OpenStudy (johnweldon1993):

Yeah I just noticed that XD, this method will still work however @ayyookyndall

OpenStudy (ayyookyndall):

Okay. So what sin(x)? Im still kinda confused.

OpenStudy (ayyookyndall):

Is there an easier way to find out?

OpenStudy (johnweldon1993):

Yeah true, probably nothing you've seen before Lets look at it the easier way

OpenStudy (anonymous):

(one leg)^2 + (the other leg)^2 = (hypotenuse)^2

OpenStudy (johnweldon1993):

 so since we know those 2 angles are equal...and the sum of all the angles in the triangle add to 180...we can solve for 'x'

OpenStudy (ayyookyndall):

Okay so the x are 21. Do I add it up with 38?

OpenStudy (johnweldon1993):

Noooo no I merely put 'x' to show that those 2 sides are equal...nothing to do with the angles...sorry So we have the angle we know..which is 38....and we know that the other 2 angles are the same so if that angle is called 'a' we would have 2a So if we add 2a + 38 we know that has to = 180...so \[\large 2a + 38 = 180\] and solve for 'x' *which will be your 'x' in your question

OpenStudy (johnweldon1993):

ugh, solve for 'a' sorry for the typo

OpenStudy (ayyookyndall):

Wow, I feel so dumb. Whats the formula to help solve for this?

OpenStudy (ayyookyndall):

Oh the answer is 70. But is there a formula that goes with this? @johnweldon1993

OpenStudy (zarkon):

70 is not correct

OpenStudy (ayyookyndall):

Really? Are you sure? @Zarkon

OpenStudy (zarkon):

you are very close

OpenStudy (ayyookyndall):

If a triangle adds up to 180 180-38= 142 142/2 = 70

OpenStudy (zarkon):

142/2=71

OpenStudy (ayyookyndall):

Oh. Gosh. Now I feel dumber. Anyways is there a formula to go with this?

OpenStudy (zarkon):

just the equation you used